OpenStudy (anonymous):

SIMPLIFY 9X - (2X +8)

OpenStudy (skullpatrol):

Any ideas where you should start?

OpenStudy (anonymous):

same is 9x + (-1)(2x+8)

OpenStudy (anonymous):

with(2x + 8)

OpenStudy (skullpatrol):

Yes.

OpenStudy (anonymous):

which gives you 10 wh

OpenStudy (anonymous):

wich would be 10x

OpenStudy (skullpatrol):

Are you saying 2x + 8 = 10?

OpenStudy (anonymous):

yeah

OpenStudy (skullpatrol):

2 + 8 = 10

OpenStudy (skullpatrol):

So you can't really do anything with what is inside (2x + 8) because one of them has an x and the other does not, understand?

OpenStudy (skullpatrol):

Use the distributive property on (-1)(2x+8) =-2x-8

OpenStudy (skullpatrol):

Now simplify 9x -2x-8

OpenStudy (skullpatrol):

(9-2)x-8=7x-8 is the answer.